OpenStudy (kimberly101):

In 3 basketball games Karina has averaged 12 points each game. In the first game she scored 11 points, in the second she scored 17 points, and in the third game she scored x points. From the given information, what is the average of 11, 17 and x?

OpenStudy (anonymous):

To find the average, you take the SUM of ALL values and DIVIDE it by the total NUMBER of values so we set up a equation based on the rules above to see how we can solve this ( 11 + 17 + x ) / ( 3 ) = 12 can you solve this?

OpenStudy (kimberly101):

is it 11+17+8 =36 =36/3 =12

OpenStudy (goformit100):